You are considering a project with an initial investment of $20 million and annual cash flow (before interest and taxes) of $5,000,000. The project’s cash flow is expected to continue forever. The tax rate is 34%, the firm’s unlevered cost of equity is 18% and its after-tax cost of debt is 6.60%. The only side-effect from the use of debt that you are concerned about is related to the tax shield.

  1. If the project were to be financed with 100% equity, would you accept the project? Why or why not?
  2. If the project were to be financed with $10 million in perpetual debt and the rest with equity, use the APV method to help you decide whether to accept the project or not. Does your decision change from part (a)?
  3. Redo part (b) by using the FTE approach.
  4. Redo part (b) by using the WACC approach.
Answer #1

a.

If the project to be financed by 100% equity

Initial Investment = $20 million

Annual Cash Flow = $5000000

Unlevered cost of Equity = 18%

Now Calculating Net Present Value (NPV)

NPV = -20000000 + 5000000/0.18 = $7777777.8

SInce net present Value is positive therfore we will accept the project.

b.

As per Adjusted Present Value Method (APV)

APV = Unlevered Firm Vale + Net effect of debt

Unlevered Firm Vale = NPV = $7777777.8

Net effect of debt = Tax Rate * Debt Amount = 0.34*10000000 = $3400000

APV = 7777777.8 + 3400000 =$11177777.8

Since APV is positive we will accept the project.

Yes, the decision will change as now the second option is preferred.

c.

Using Free Cash Flow method

Annual Cash Flow = $5 million

Interest Expense = 0.066*10 = $0.66 million

Net Cash Flow = $(5-0.66) million = $ 4.34 million

Tax Rate = 34%

Tax Amount = $1.4756 million

Free Cash Flow =( $5-$1.4756)million = $3.5244 million

NPV = (-20 + 3.5244/0.18 ) million= -$0.42 million

d.

Debt = $10 million

Equity = $10 million

WACC = (D/D+E)* After tax Cost of Debt + (E/D+E)* Cost of Equity

WACC = (10/20)*0.066 + (10/20)*0.18 = 12.3%

Now Calculating NPV

NPV = -$20 million + $5 million/0.123 = $20.65 million